The temperature dropped 15 degrees celsius in the last 30 days. If the rate of temperature drop 
remains the same, how many degrees will the temperature drop in the next ten days? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Temperature drop in 30 days = 150C Temperature drop in 1 day = 15 / 30
= (1 / 2)0 C
So, temperature drop in next 10 days = (1 / 2) × 10
= 50C
∴The temperature drop in the next 10 days will be 50 C

Thanks